### v1

# OpenLKA: An Open Dataset of Lane Keeping Assist from Recent Car Models under Real-world Driving Conditions

Abstract
Lane Keeping Assist (LKA) is widely adopted in modern vehicles, yet its real-world performance remains underexplored due to proprietary systems and limited data access. This paper presents OpenLKA, the first open, large-scale dataset for LKA evaluation and improvement. It includes 400 hours of driving data from 50+ production vehicle models, collected through extensive road testing in Tampa, Florida and global contributions from the Comma.ai driving community. The dataset spans a wide range of challenging scenarios, including complex road geometries, degraded lane markings, adverse weather, lighting conditions and surrounding traffic. The dataset is multimodal, comprising: i) full CAN bus streams, decoded using custom reverse-engineered DBC files to extract key LKA events (e.g., system disengagements, lane detection failures); ii) synchronized high-resolution dash-cam video; iii) real-time outputs from Openpilot, providing accurate estimates of road curvature and lane positioning; iv) enhanced scene annotations generated by Vision Language Models, describing lane visibility, pavement quality, weather, lighting, and traffic conditions. By integrating vehicle-internal signals with high-fidelity perception and rich semantic context, OpenLKA provides a comprehensive platform for benchmarking the real-world performance of production LKA systems, identifying safety-critical operational scenarios, and assessing the readiness of current road infrastructure for autonomous driving. The dataset is publicly available at: https://github.com/OpenLKA/OpenLKA.
Index Terms:
Lane Keeping Assist, Autonomous Driving Dataset, CAN–Video Sync Logging, Automatic dataset annotation, Real-World Driving Data
Figure 1: OpenLKA Dataset Overview
I Introduction
Road-departure crashes account for a substantial share of traffic fatalities, and inadvertent lane-keeping errors are recognised as their primary cause [1]. To mitigate this risk, automakers have accelerated the roll-out of Advanced Driver Assistance Systems (ADAS) [2, 3, 4]. Lane-Keeping Assist (LKA) has progressed from a premium option to a default feature, offering continuous steering support intended to keep vehicles centred in their lanes [5, 6, 7]. Yet field reports show that production LKA stacks can drift or disengage on sharp curves, worn lane markings, or in adverse weather [8]; drivers therefore remain uncertain about when the system will intervene or relinquish control [9]. At the same time, infrastructure agencies lack quantitative guidance on lane-marking contrast or geometric design that accommodates machine vision [10], and researchers are hampered by limited access to real-world failure logs—information manufacturers seldom share [11].
While a few transportation departments have begun examining LKA behaviour in specific settings [12], such studies involve only a handful of vehicles and roadway conditions. The research community therefore still lacks a large-scale, multimodal corpus centred on LKA—one that spans multiple makes, climates, and driving styles—to support systematic benchmarking and infrastructure evaluation.
To bridge this gap we introduce OpenLKA, an openly available dataset that fuses full CAN messages, high-resolution dash-cam video and lane-detection outputs from production vehicles; the corpus combines a controlled campaign—19 vehicle models driven by three professional drivers for 2 986 LKA-engaged trips (19 380 min) in Tampa, Florida—with 775 community-contributed drives (3 964 min) from 52 additional models that interleave LKA operation and purely human control, thereby yielding in total 61 car models, 95 drivers, 3 761 drives and 400 hours of synchronised CAN–video data, a scale that not only exposes brand- and environment-dependent LKA behaviour but also enables direct human-versus-automation comparisons and rich analyses of driver–LKA interaction in everyday traffic, as summarised in Figure 2.
The contributions of this work are as follows:
  1. (i)
    A first-of-its-kind open dataset for real-world LKA analysis: OpenLKA provides 400 hours of multimodal driving data from 50+ production vehicles models across various car makes, collected under diverse and challenging real-world conditions, including adverse weather, complex road geometries, and degraded lane markings.
  2. (ii)
    Reverse-engineering black-box OEM signals for LKA introspection: We develop and apply custom reverse-engineered DBC files to decode internal LKA-related signals from raw CAN bus data—such as system disengagements, lane detection failures, and human driver steering takeover—which are otherwise inaccessible due to the proprietary nature of commercial vehicle systems.
  3. (iii)
    A unified LKA data collection pipeline and ecosystem: We introduce a scalable data collection framework that captures full CAN bus streams, synchronized dash-cam video, and real-time outputs from Openpilot, enabling scalable estimation of road curvature and lane positioning that would otherwise require costly LIDAR or manual surveys. This pipeline further powers an automated ecosystem that continuously ingests new LKA and human steering data from global Comma.ai community drivers operating in diverse locations and road conditions.
  4. (iv)
    VLM-based rich contextual annotation: Each driving segment is automatically labeled using Vision Language Models (VLM), with proper prompt engineering, the dataset captures accurate and detailed environmental factors including lane visibility, pavement quality, weather, lighting, and traffic.
The remainder of the paper is organised as follows. Section II surveys related datasets and LKA field studies. Section III describes the hardware and software pipeline used to collect and synchronise CAN and video data. Section IV presents dataset structure and statistics. Section V explains the annotation and quality-assurance procedures. Section VI outlines research directions that OpenLKA enables, and Section VII concludes with future work.
II Related Work
LKA systems, a cornerstone of ADAS, utilize computer vision (CV) techniques to detect lane lines and employ control algorithms to maintain vehicles within lane boundaries, significantly enhancing driving safety [13]. These systems typically rely on convolutional neural networks (CNNs) to process camera inputs, achieving lane detection accuracies exceeding 97% on benchmark datasets like TuSimple [14]. Despite advancements in CV, lane detection remains vulnerable to long-tail scenarios such as adverse weather conditions, poor lighting, or unconventional lane markings, issues exacerbated by imbalanced training datasets that do not represent diverse global conditions [15]. Moreover, LKA disengagements are not solely attributable to lane detection failures; system designs encompass multiple factors, including sensor fusion and driver interaction, underscoring the necessity of empirical evaluations of market-ready products [16]. Perception failures can propagate to downstream planning modules, where suboptimal design may impair effective risk mitigation, and control systems occasionally fail to maintain consistent lane centering, posing safety hazards [17]. Given the proprietary, black-box nature of LKA systems, their evaluation is particularly crucial through methods such as AI explainability, which can elucidate perception failures and guide infrastructure improvements (e.g., optimized lane markings) to enhance safety and mitigate lane departure accidents—a leading cause of road fatalities [18]. This synergy between technological capabilities and infrastructure quality highlights the ongoing need for robust, transparent, and comprehensive LKA system research.
Currently, the design and evaluation of ADAS, LKA, and autonomous driving technologies are critically hindered by the lack of comprehensive, multimodal datasets that capture diverse real-world conditions [3, 2]. Existing datasets, such as KITTI [19] and BDD100K [20], primarily focus on sensor and video data but lack crucial CAN bus integration, limiting their applicability in analyzing detailed vehicle dynamics and internal system states critical to understanding LKA behavior. Moreover, many of these datasets are narrowly scoped toward fully autonomous driving scenarios, for example NuScenes  [21], neglecting nuanced human-LKA interactions and failing to represent diverse global driving conditions adequately. This shortcoming restricts researchers from conducting comparative studies between human-driven and automation-driven behaviors, essential for advancing safety and reliability. Additionally, most existing datasets are limited in environmental variability, insufficiently capturing adverse weather, degraded road infrastructure, and complex lighting conditions that are prevalent in real-world driving scenarios. Such limitations impede the development and validation of robust LKA systems capable of handling real-world operational complexities.
Figure 2: Vehicle Distribution in the OpenLKA Dataset
III Experiments, Data Collection and Processing
OpenLKA addresses the critical need for real-world, multimodal, scalable datasets for LKA research. Key design principles include extensibility, openness, multimodal richness and contextual diversity. Unlike existing datasets, OpenLKA integrates synchronized CAN bus, dashcam video, openpilot outputs, and contextual scene annotations, making it uniquely comprehensive.
III-A Experiment Hardware Setup
To evaluate the real-world performance of LKA systems in consumer vehicles [22, 23], we utilized the Comma Three-X [24], an aftermarket self-driving development kit. The Comma Three-X was selected for its integrated high-definition front-view camera and access to the Openpilot ecosystem, which facilitates synchronized data collection. The device connects to the vehicle’s internal Controller Area Network (CAN) bus via the physical On-Board Diagnostics (OBD-II) port, enabling the extraction of high-frequency control signals through reverse engineering, as standard OBD-II protocols do not provide such detailed data.
The Comma Three-X was mounted in the center of the vehicle’s interior windshield (Fig. 3) after rigorous calibration of its camera to ensure accurate lane deviation measurements. Its internal sensors operate at a frequency of 100 Hz, and its state-of-the-art lane detection algorithm, which employs an end-to-end approach based on computer vision and machine learning, identifies lane lines in real time. However, this algorithm serves as a reference and does not reflect the performance of the vehicle’s native LKA system. By synchronizing CAN data with video footage, we obtained a comprehensive view of the vehicle’s mechanical responses in relation to its driving environment.
To eliminate interference with the vehicle’s native LKA system, the Comma Three-X was operated in dash-cam mode, where it passively collects CAN bus data without exerting control. We confirmed the absence of interference through electrical load testing and monitoring of the vehicle’s native system status, ensuring the integrity of the stock LKA system during data collection.
Figure 3: Comma Three-X dash cam setup for LKA data collection, mounted at the center of the vehicle’s windshield.
III-B Selection of Testing Vehicles and Scenarios
TABLE I: Summary of Testing Scenarios
Variable Conditions Tested
Weather Clear, Rain, High Rain, Storm, Thunderstorms, Fog, Extreme Heat
Lighting Daytime, Nighttime, Dawn, Dusk, Sun glare, Street lights, Tunnel transitions
Pavement Quality Smooth asphalt, Rough surfaces, Potholes, Wet surfaces, Gravel, Repaired patches
Lane Markings Well-defined, Faded, Obscured, Absent, Reflective markers, Different Color Combinations, Transitions
Road Geometry Straight, Sharp curves, Grades, Intersections, Construction zones, Merges/splits, Roundabouts, Tunnels
Traffic Free-flow, Light, Moderate, Heavy Congested, Stop-and-go, Traffic Jam
Vehicle Speeds Low (<<<20 mph), Medium (20–45 mph), High (>>>45 mph), Variable speeds
Figure 4: Different of test vehicles across manufacturers.
Our experimental design aimed to comprehensively assess LKA performance under diverse real-world driving conditions. We conducted tests using a fleet of rental cars sourced from Tampa International Airport, featuring models from major manufacturers such as Honda, Toyota, Tesla, Ford, Kia, Hyundai and so on (Fig. 4). This selection ensured broad coverage of LKA implementations across different automakers.
Testing was primarily conducted in Tampa, Florida, and surrounding areas, leveraging the region’s subtropical climate to evaluate LKA systems under challenging conditions, such as heavy rain, high humidity, sharp curves, unclear lane markings, and tunnels, summarized in Table I. Routes were systematically selected to include complex road segments (Fig. 5), and tests were performed across various times of day and weather conditions to ensure comprehensive coverage (Fig. 6). Additionally, we incorporated data from a community dataset, which, despite its randomized collection nature, captured diverse weather, traffic, and pavement conditions from multiple regions, further enhancing the dataset’s diversity.
Figure 5: LKA data collection routes in Tampa, Florida, covering I-275, I-75, and two-lane rural road.
Figure 6: Examples of testing scenarios.
All tests were conducted by our team, consisting of experienced drivers with extensive driving backgrounds. We adhered to a protocol of utilizing the LKA system whenever feasible, provided it did not result in significant vehicle deviation or disengagement, ensuring consistent evaluation of the system’s performance under varied conditions.
III-C CAN Log Data Acquisition and Reverse Engineering
CAN logs were acquired using the Comma Three-X by accessing the vehicle’s internal CAN bus through the OBD-II port. The logs were recorded in two formats: qlog files (10 Hz, for low-frequency state monitoring) and rlog files (100 Hz, for high-frequency dynamic analysis). We employed open-source DBC files from OpenDBC to decode CAN messages, extracting variables such as clock, LKA/ACC status, pose angles, steering metrics, lane detection, lane offset, object detection, dynamics, and jerk. To ensure accuracy, we manually validated DBC files for Toyota, Kia, and Ford models via signal injection and observation, while Tesla models (e.g., Model 3, Model X) required reverse engineering for precise control signals. Signal availability per vehicle model is detailed in Table II.
Figure 7: Two examples from OpenLKA, the left are the video frames captured by OpenPilot’s dash-cam; the middle are the lane position estimate by the OpenPilot’s end-to-end model; The right are the lane detections results from CAN.
TABLE II: Signal Availability Across Vehicle Models
Model Clk ACC Pose Steer Lane Det. Obj. Det. Dyn. Jerk
Volkswagen Y Y Y Y Y Y Y Y
Ford Y Y Y Y N Y Y N
Toyota Y Y Y Y Y N Y N
Honda Y Y Y Y Y Y Y Y
Tesla Y Y Y Y (RE) Y (RE) Y Y Y
Kia Y N Y Y N Y Y N
Y: Yes, N: No, RE: Reverse Engineered.
III-D Scalable Lane Position Estimation via Openpilot
We integrate Openpilot, a state-of-the-art open-source advanced driver assistance system, into our dash-cam setup to capture high-fidelity perception outputs alongside front-facing video. Openpilot’s robust vision-based model estimates critical lane-keeping parameters in real time, including the vehicle’s lateral offset from lane markings and the curvature of the road.
These estimations serve as effective proxies for ground truth in our dataset. While not perfectly precise, they offer a practical and scalable alternative to traditional methods such as LiDAR-based lane detection or manual road geometry surveys, which are significantly more costly and labor-intensive. Openpilot thus enables efficient collection of supervisory signals across large-scale, real-world driving data.
Figure 8: In (a) and (b), the left images are the different time driving in similar place captured by Openpilot’s dash-cam, and the middle and right part are both from Openpilot’s output.
By running Openpilot in a passive, observational mode, we preserve the native behavior of the stock LKA system while simultaneously generating consistent, model-based perception outputs. This dual-view setup enables side-by-side comparisons between proprietary systems and a common reference model, supporting fine-grained analysis of lane centering and lateral control performance. An example is shown in Fig. 8, where the vehicle’s lateral deviation from the lane center is estimated and plotted. Manual inspection confirms that the temporal gradient of this deviation closely matches the vehicle’s measured lateral velocity, suggesting strong internal consistency and high precision in Openpilot’s perception outputs. This empirical alignment supports the use of Openpilot as a practical proxy for ground-truth lane position in large-scale observational studies.
III-E VLM-Based Rich Contextual Annotations
TABLE III: Description of Annotated Labels
Category 0 1 2 3 4 5 6
Lane Mark. Worn out Partially worn Intact - - - -
Weather Rain Dusty Clear Snow - - -
Lighting Dark Artificial light Normal light - - - -
Traffic Traffic jam Stop-and-go Congested Free flow - - -
Road Cond. Large potholes Small cracks Smooth - - - -
Driving Reckless Major anomaly Minor anomaly Normal - - -
Pedestrian None Few Moderate Crowded - - -
Road Type Two-lane Three-lane Four-lane Multi-lane Roundabout Intersection -
Scenarios None Oncoming vehicle Emergency vehicle Construction Obstacle Animal Accident
Surr. Veh. None Low density Moderate density High density - - -
To provide richer contextual information for detailed analysis of LKA scenarios descriptions, we annotated the OpenLKA video data using VLM, specifically GPT-4o [25]. The video footage was segmented into individual frames, each analyzed together with its immediate preceding and subsequent frames. This method reduces visual obstructions (e.g., windshield wipers, water droplets) and enhances temporal context, leading to more robust and accurate annotations (see Fig. 9).
We developed a domain-specific annotation prompt tailored explicitly for transportation scenarios within the OpenLKA dataset. The prompt utilized System Prompting, Role Prompting and few-shot prompting techniques, instructing the VLM to act as an experienced transportation domain expert [26]. Additionally, the prompt incorporated the Chain-of-Thought (CoT) methodology, promoting systematic, step-by-step reasoning to improve consistency and reduce annotation hallucinations [27]. Few-shot examples were also included to clearly illustrate the intended annotation process. The VLM classified each frame into predefined environmental categories (as shown in Table III). Annotation accuracy was validated by manually labeling a representative subset of 2,000 images, selected via stratified random sampling across all categories. Comparing manual annotations with GPT-4o’s outputs revealed an accuracy exceeding 85%. Notably, the VLM exhibited a conservative labeling tendency, frequently identifying scenarios as more severe than manual annotations, a beneficial characteristic for safety-critical transportation system design.
Figure 9: Annotation pipeline for video frames using VLM.
III-F Ecosystem with Community-Driven Data Streaming
OpenLKA uniquely incorporates continuous updates from the global comma.ai self-driving community, marking it the largest LKA dataset to leverage community-sourced data streams. Our open-source pipeline automatically converts community-collected CAN/video data into OpenLKA-compatible formats. Users worldwide can seamlessly contribute data, and our processing tools (CAN decoding, synchronization scripts, VLM annotation methods) are fully open-sourced. We anticipate OpenLKA evolving into one of the most diverse LKA datasets globally, empowering researchers to create customized datasets using our publicly available tooling.
IV Dataset Overview
TABLE IV: Overview of OpenLKA Dataset Contents
Category Component Description
[…]
CAN & Openpilot System Status Activation states for Openpilot, adaptive cruise control, and lane-keeping systems.
Vehicle Dynamics Vehicle motion data including speed, acceleration, steering, and body orientation.
Perception Outputs Environmental awareness data such as lane visibility, lane departure risks, and lead vehicle detection.
Planning Outputs Path planning data including desired curvature and predicted trajectories.
VLM Annotations annotation.json Environmental factors: weather, traffic congestion, lighting, road markings.
Other DBC Files Reverse-engineered CAN DBC files for various vehicle models.
Prompt Template Chain-of-Thought prompt for VLM annotation.
The OpenLKA dataset was developed to address the need for a comprehensive, multimodal resource tailored for LKA system research, offering a diverse and parameter-rich foundation for analyzing real-world LKA performance and human-LKA interactions. Spanning 389.07 hours across 1,166 unique driving routes, the dataset encompasses data from 50+ vehicle models and nearly 100 drivers, organized in a structured car model, dongle id, route id format to distinguish between models, drivers, and trips. In the first, self-collected LKA data, was gathered starting in May 2024 by three drivers using 19 market-leading vehicle models rented from Tampa International Airport, covering a wide range of brands prevalent in the U.S. market, resulting in 323 hours (83% of the total duration) of driving data across Florida’s diverse weather, lighting, road, and traffic conditions; the second component, sourced from the Comma autonomous driving community, comprises 66.1 hours (17%) of human driving data from 92 drivers across 50+ vehicle models worldwide, collected as of April 2025, providing a valuable baseline for comparing human driving styles with LKA performance and exploring human-LKA interaction dynamics, as illustrated by the global distribution in Figure 10 [24]. Additionally, this part of the dataset covers more than 30 states, provinces, and territories in over 80 cities around the world. This greatly enriches the dataset with driving scenarios such as snow, dust storms, and frosty roads, as well as driving scenarios in mountainous areas, dense cities, and deserts.
Figure 10: Global Distribution of Community Driving Dataset
The dataset’s diversity is evident in its composition, the most represented model, HYUNDAI IONIQ 5 2022, contributes 14.61% of the total duration (56.85 hours). The dataset includes synchronized multimodal streams, such as videos at multiple resolutions, vehicle control data via CAN, Openpilot’s end-to-end model outputs, and VLM-annotated environmental factors, with a detailed breakdown provided in Table IV. Privacy is ensured by blurring faces and license plates in all videos to protect driver identities.
Figure 11: LKA performance overview under normal conditions
Table V compares OpenLKA with existing autonomous driving datasets, underscoring its unique strengths: comprehensive environmental annotations covering weather, visibility, traffic, and infrastructure; synchronized video and CAN data for precise temporal alignment; rich CAN signals paired with Openpilot’s baseline outputs; and VLM-augmented labels that enhance contextual understanding, positioning OpenLKA as a uniquely diverse and parameter-rich resource for advancing LKA system development and dataset-driven research in autonomous driving.
TABLE V: Comparison of OpenLKA and Existing Datasets
Dataset Data Env LLM Time
Vis CAN LiD Wthr Vis Traf Infra Ann (hrs)
OpenLKA 1 1 0 1 1 1 1 1 390+
HDD 1 0 0 1 1 1 1 0 104
BDD-X 1 0 0 1 1 1 1 1 77
WoodScape 1 0 0 1 1 1 1 0 N/A
DAWN 1 0 0 1 1 1 1 0 N/A
KITTI 1 0 1 0* 0 0 0 0 1.5
BDD100K 1 0 0 1 1 1 1 0 1.2K
Comma2k19 1 1 0 0 0 0 0 0 33
Cityscapes 1 0 0 1 1 1 1 0 N/A
ApolloScape 1 0 1 1 1 1 1 0 N/A
nuScenes 1 0 1 1 1 1 1 0 1.2K+
Waymo 1 0 1 1 1 1 1 0 570+
*Limited weather conditions in KITTI dataset.
V Dataset Applications
The OpenLKA dataset, with its rich multimodal and diverse real-world driving data, offers a robust foundation for advancing research in intelligent transportation systems (ITS) and autonomous driving (AD). Below, we outline several potential applications that leverage the dataset’s unique characteristics.
  1. 1.
    LKA benchmarking and real-world evaluation: This dataset contains real-world data of LKA systems of different vehicle makes in different scenarios, which we can use to quantify the level of LKA in different car makes. In addition, we can also use it to quantify the most challenging scenarios of LKA and analyze the causes, thus providing guidance for LKA development and traffic safety prevention.
  2. 2.
    Enhanced LKA Design Through multimodal Data and Foundation Models: The OpenLKA dataset provides a comprehensive foundation for designing more robust LKA systems by offering diverse variables related to lane line detection and vehicle control dynamics. We also have end-to-end model outputs from OpenPilot, serving as a valuable baseline for comparative analysis. To demonstrate the dataset’s utility for advanced driving automation research, we conducted experiments with multimodal language models as shown in Fig. 12. We sampled representative driving scenarios from OpenLKA, preparing synchronized video frames and CAN bus data as inputs, with human drivers’ subsequent steering angles as labels. Through Supervised Fine-Tuning (SFT) using ChatGPT-4o’s API on these multimodal inputs, we observed significant improvements in steering prediction accuracy. These results highlight the considerable potential of leveraging multimodal foundation models for automated driving decision-making based on OpenLKA.
  3. 3.
    Infrastructure Readiness Assessment and Roadway Safety Planning: OpenLKA enables transportation infrastructure stakeholders to evaluate how current road conditions interact with LKA system performance. By identifying segments where LKA systems frequently disengage or exhibit poor lane-keeping—especially under conditions such as degraded lane markings, sharp curves, or inconsistent geometry—stakeholders can assess the roadway’s readiness for assisted driving technologies. The dataset facilitates targeted interventions, such as pavement re-striping, lane marking restoration, or geometric redesign, to reduce LKA failure risk and improve overall lane-keeping safety at the infrastructure level.
  4. 4.
    Analysis of LKA Failure to Driver Takeover Dynamics: When LKA systems fail or exhibit significant lane deviation, drivers often have little to no time to react, posing safety risks. The OpenLKA dataset provides a comprehensive record of the entire process from LKA failure or large lane deviation to driver takeover, capturing multimodal signals such as CAN data (e.g., steering angle), video footage, and VLM-annotated environmental contexts (e.g., faded markings, sharp curves). This enables researchers to analyze LKA failure modes, quantify the time window from failure to takeover, and model driver reaction patterns under diverse conditions [28]. Such insights can inform the development of proactive warning systems that predict LKA instability and alert drivers in advance, enhancing overall safety.
Figure 12: Steering angle prediction using GPT-4o: The original model predicts 6.85°, while fine-tuning with OpenLKA data yields 7.52°, matching the human driver’s actual action. (7.52°).
VI. Conclusion
The OpenLKA dataset pioneers a uniquely comprehensive, multimodal, and real-world foundation for advancing LKA system research. Its primary strength lies in its integration of synchronized CAN bus, high-definition dashcam videos, detailed Openpilot outputs for precise lane-keeping metrics, and extensive contextual annotations via VLM. Furthermore, OpenLKA uniquely leverages continuous data contributions from a large, global driving community, ensuring ongoing dataset expansion and diversity.
However, it lacks data on trailing vehicles, which is critical for comprehensive traffic studies. Future work will integrate radar and LiDAR to enhance data collection, and leverage the dataset to evaluate LKA performance in real-world scenarios, propose improvements for in-vehicle systems and roadside infrastructure, and develop human-centered warning systems for smoother ADAS takeovers.